undeadknight26 (undeadknight26):

Identify the factors of x^2 + 36y^2. Prime (x + 6y)(x - 6y) (x + 6y)(x + 6y) (x - 6y)(x - 6y)

undeadknight26 (undeadknight26):

@SolomonZelman @surjithayer

OpenStudy (mathmale):

If you were really stuck, you could try multiplying out the factors given in B, C and D. Try one of those. Do the results equal x^2 + 36y^2 ?

undeadknight26 (undeadknight26):

@linda3 @linn99123 @agent0smith

OpenStudy (linn99123):

o.0

OpenStudy (mathmale):

You are welcome to contact anyone you want for help, but the suggestions I have already given you should be all you really need to know to select the correct answer to this question. Have you tried the multiplication I've suggested? Isn't your objective to learn how to solve problems by yourself?

undeadknight26 (undeadknight26):

Hey sir...I know how to do these things on my own...im just having problems with this one. I also have 10 more questions i have to do in 10 minutes and these do not take me 1 minutes per question....

undeadknight26 (undeadknight26):

So can some one explain it to me through numbers lolz?

OpenStudy (agent0smith):

Do as @mathmale did. I'll start the first for you... (x + 6y)(x - 6y) = x^2 + 6xy - 6xy - 36y^2 = (x + 6y)(x + 6y) = (x - 6y)(x - 6y) =

undeadknight26 (undeadknight26):

thanks...

OpenStudy (mathmale):

If you really have ten more questions to do in ten minutes, move on, please, and return to this question if you have any time left. @agent0smith: t y very much.

OpenStudy (agent0smith):

Are you not able to finish multiplying/expanding those out...? it will take seconds.